The table below gives the number of hours spent unsupervised each day as well as the overall grade averages for seven randomly selected middle school students. Using this data, consider the equation of the regression line, yˆ=b0+b1xy^=b0+b1x, for predicting the overall grade average for a middle school student based on the number of hours spent unsupervised each day. Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant.

Hours Unsupervised 1.51.5 2.52.5 33 3.53.5 4.54.5 55 66
Overall Grades 8989 8686 7777 7474 7373 7272 7171

Step 1 of 6:

Find the estimated slope. Round your answer to three decimal places.

Step 2 of 6:

Find the estimated y-intercept. Round your answer to three decimal places.

Step 3 of 6:

According to the estimated linear model, if the value of the independent variable is increased by one unit, then the change in the dependent variable yˆy^ is given by

Step 5 of 6:

Find the error prediction when x=5x=5. Round your answer to three decimal places.

Step 6 of 6:

Find the value of the coefficient of determination. Round your answer to three decimal places.
Answer #1
X Y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
1.5 89 4.9030612 133.898 -25.6224
2.5 86 1.4744898 73.46939 -10.4082
3 77 0.5102041 0.183673 0.306122
3.5 74 0.0459184 11.7551 0.734694
4.5 73 0.6173469 19.61224 -3.47959
5 72 1.6530612 29.46939 -6.97959
6 71 5.2244898 41.32653 -14.6939
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 26 542 14.42857 309.7143 -60.1429
mean 3.714285714 77.4285714 SSxx SSyy SSxy

sample size ,   n =   7      
here, x̅ =   3.714285714       ȳ =   77.42857143
              
SSxx =    Σ(x-x̅)² =    14.42857143      
SSxy=   Σ(x-x̅)(y-ȳ) =   -60.14285714      

correlation coefficient ,    r = Sxy/√(Sx.Sy) =   -0.8997
---------------------

correlation hypothesis test      
Ho:   ρ = 0  
Ha:   ρ < 0  
n=   7  

df=n-2 = 5


alpha,α =    0.05  
correlation , r=   -0.8997  
t-test statistic =    t = r*√(n-2)/√(1-r²) =    -4.6085
critical t-value = -2.0150
p-value =    0.0029  
since,p-value<α , reject Ho

hence, correlation is significant .

--------------------------------------------------------------------------------------------------------------

1)

SSxx =    Σ(x-x̅)² =    14.42857143
SSxy=   Σ(x-x̅)(y-ȳ) =   -60.14285714

slope ,    ß1 = SSxy/SSxx =   -4.168

2)

intercept,   ß0 = y̅-ß1* x̄ =   92.911

3)

so, regression line is   Ŷ =   92.9109   +   -4.1683   *x

According to the estimated linear model, if the value of the independent variable is increased by one unit, then the change in the dependent variable yˆ is given by decrease by 4.1683

5) when x=5

then

so, regression line is   Ŷ =   92.9109   +   -4.1683*5 = 72.069

error= observed-predicted value= 72-72.069 = - 0.069

6)

value of the coefficient of determination = r² = -0.8997² = 0.809